# Impurity bands, line-nodes, and anomalous thermal Hall effect in Weyl superconductors

### v1

@@ Abstract
We investigate the anomalous thermal Hall effect (ATHE) in Weyl superconductors realized by the $E_{1u}$ ($p$-wave and $f$-wave) chiral superconducting order for the point group $D_{6h}$. Using the quasiclassical Eilenberger theory, we analyze the influence of the impurity scattering and the line nodal excitations on the ATHE, and compare it with the intrinsic (topological) contribution. Because the transverse response is sensitive to the slope of the density of states at the Fermi surface, the extrinsic ATHE vanishes in both the weak (Born) and strong (unitarity) scattering limits. The thermal Hall conductivity (THC) is maximal at intermediate impurity strengths when there is a large slope of the density states in the impurity bands close to the Fermi energy. Under these conditions, the extrinsic ATHE dominates the intrinsic ATHE even at low temperatures. The extrinsic ATHE is sensitive to line nodal excitations, whereas the intrinsic ATHE is not. When the line nodes in the gap involve the sign change of the order parameter, the extrinsic contribution to the THC is suppressed even though the phase space for low energy excitation is large. In contrast, if the nodes are not accompanied by such a sign change, the extrinsic ATHE is significantly enhanced. Our results form a basis for the comprehensive analysis of anomalous thermal transport in Weyl superconductors.
@@ I Introduction
Weyl superconductors (WSCs) are time-reversal symmetry broken (TRSB) superconductors whose low-energy excitations behave as Weyl quasiparticles [1]. They are realized when the superconducting condensate consists of Cooper pairs with a fixed orbital angular momentum and thus are described by a complex order-parameter, $\Delta(\bm{k})\propto(k_{x}\pm ik_{y})^{\nu}\;(\nu\in\mathbb{Z})$. On a three-dimensional Fermi surface centered on the $\Gamma$ point, the gap closing points at $k_{x}=k_{y}=0$ become Weyl nodes, which are sources and drains of Berry flux [2, 3, 4]. As a result, the near-nodal Bogoliubov quasiparticles behave as Weyl particles. It is common to both refer to such order parameters and label the corresponding ground states as chiral [5]. Generically, chiral superconductors with a three-dimensional Fermi surface are good candidates for the realization of WSCs.
@@
Among the phenomena that experimentally identify WSCs are the anomalous thermal Hall effect (ATHE), a transverse thermal current driven by a temperature gradient without an applied magnetic field, and the chiral anomaly-induced phenomena, such as the torsional chiral magnetic effect and the negative thermal magnetoresistivity by textures of the order parameters [6, 7, 8, 9, 10, 11, 12, 13]. In particular, the observation of the ATHE unambiguously identifies the chiral ground states.
@@
At the microscopic level, there are two sources for a finite thermal Hall signal in chiral superconductors: intrinsic and extrinsic. The intrinsic mechanism is due to the geometric phase from the Berry curvature and characterized by the positions of Weyl nodes in the momentum space [14, 15, 16, 17, 18, 19]. The extrinsic mechanism is due to impurities and to the transfer of the angular momentum between the condensate and the quasiparticles during scattering events [20, 21, 22, 23, 24], which contributes both to energy-dependent skew scattering of quasiparticles on impurities and the Andreev (inter-branch, electrons to holes and vice versa) scattering. The skew scattering directly couples to the temperature gradient for $|\nu|=1$ and results in the ATHE signal [20, 21, 22]. The Andreev mechanism only couples to impurities if the impurity potential is non-$s$-wave, with scattering matrix elements coupling different angular-momentum channels, and therefore appears for finite size impurities, dominating the response for $|\nu|\geq 2$ [23, 24, 25].
@@
In this paper, we address two aspects of the ATHE in chiral superconductors with point-like impurities. First, we elucidate its origin and demonstrate the relation between the extrinsic ATHE and the evolution of the density of states (DOS) in the sub-gap impurity band, which arises from the broadening of the impurity resonant states [26, 27]. In particular, the particle-hole anisotropy, which is necessary for transverse transport, sensitively depends on the scattering phase-shift at individual impurities [28, 29]. Second, we focus on the magnitude of the thermal Hall conductivity (THC) in the situation where the winding number, $\nu$, is not the same as the total angular momentum of the Cooper pairs, $l$. Such a situation often occurs when chirality coexists with additional nodes in the gap function.
@@
Our results are relevant to several candidates of WSCs. Historically, the Anderson-Brinkmann-Morel (ABM) state in 3He was the first well-established Weyl superfluid with the chiral $p$-wave pairing [30, 1, 31, 32, 33, 34, 35]. In superconducting materials, the chiral ground states over three-dimensional Fermi surfaces were proposed in a number of materials including URu2Si2, UPt3, U1-xThxBe13 and SrPtAs. [36, 37, 38, 15, 39, 40, 41, 42, 43, 44, 45, 46, 47, 16, 48, 49, 50, 51, 52, 53, 54, 55] The chiral $d$-wave pairing, $\Delta(\bm{k})\propto k_{z}(k_{x}\pm ik_{y})$, was proposed in URu2Si2 and explains the giant Nernst effect above the critical temperature due to preformed chiral Cooper pairs. [39, 40] Uranium compounds UPt3 and U1-xThxBe13 show spin-triplet superconductivity and multiple superconducting phases as a function of tuning parameters. [56, 57, 58, 59, 60] While the exact order parameter symmetry in these materials is still a subject of debate, TRSB was observed in the so-called B-phase, suggesting chiral superconducting nature [43, 61]. In general, ferromagnetic superconductors show a complex “nonunitary” order parameter and support Weyl nodes on the three-dimensional Fermi surface [62, 63, 64, 65]. Consequently, ferromagnetic UCoGe, URhGe and UGe2 are also good candidates for the realization of WSCs [66, 67, 68, 69].
@@
We employ the quasiclassical method which is a hierarchical expansion in $1/(k_{\rm F}\xi_{0})\sim T_{\rm c}/T_{\rm F}\ll 1$, where $k_{\rm F}$ is the Fermi momentum, $\xi_{0}=v_{\rm F}/(2\pi T_{\rm c})$ is the superconducting coherence length, and $T_{\rm c}$ and $T_{\rm F}$ are the superconducting transition and the Fermi temperature respectively. Throughout the paper, we set $k_{\rm B}=\hbar=1$ for simplicity. In this language, the extrinsic contribution is of leading order, while the intrinsic mechanism is smaller by $1/(k_{\rm F}\xi_{0})$. Hence, understanding the extrinsic contribution is essential for correctly interpreting future experiments measuring the ATHE. However, while the intrinsic ATHE has been widely investigated in the context of topological material science [70, 71, 72, 73, 74, 7, 75, 16, 17, 18, 19], the studies of the extrinsic ATHE have been very limited. Our work thus fills this gap for Weyl superconductors.
@@
The rest of this paper is organized as follows. We introduce the model of WSCs in Sec. II and present the expression of the THC in Sec. III. In Sec. IV, we review the quasiclassical transport (Eilenberger) theory. In Sec. IV.3, the nonequilibrium Green function is calculated. The result of the low temperature analysis is given in Sec. V. In Sec. VI, we use the result of the low temperature analysis to connect the extrinsic ATHE and the formation of the impurity band. In Sec. VII, the influence of line nodal excitations on the ATHE is discussed. Section VIII is devoted to a summary and conclusion. Detailed derivation of several more complex equations in the main text is given in the Appendices.
@@
Figure 1: The superconducting gap structures in the $E_{1u}$ chiral superconducting states considered in the main text [see Eqs. (1) and (2)]. The left and right panels illustrate the superconducting gap structure of the chiral $p$-wave and $E_{1u}$ chiral $f$-wave states, respectively. The color of the superconducting gap represents the phase of the complex order parameter. The grey sphere indicates the underlying Fermi sphere.
II Model
In this paper, we consider the chiral superconducting order for the point group $D_{6h}$ and short-range impurities approximated by a delta-function potential. Table. 1 summarizes the irreducible representations and their basis functions exhibiting the chiral superconducting order. [76] Among them, the skew scattering does not couple to the $E_{1g}$, $E_{2g}$ and $E_{2u}$ chiral superconducting order with the $\delta$-function type potential. [23, 24] When impurities have finite size, the Andreev scattering couples to these chiral superconducting orders and contributes to the ATHE. Thus, the extrinsic mechanism is negligible in the $E_{1g}$, $E_{2g}$ and $E_{2u}$ chiral states when the impurity potential is short-range.
The $E_{1u}$ chiral superconducting order has the $p$-wave and $f$-wave pairings. As shown below, these pairing states couple to the skew scattering and cause the extrinsic ATHE. Hence, we consider the $p$-wave and $f$-wave pairings in the $E_{1u}$ state on a spherical Fermi surface.
The $p$-wave and $f$-wave chiral pairings in the $E_{1u}$ state are expressed by,
with the normalized momentum $\hat{\bm{k}}={\bm{k}}/k_{\rm F}$. Here, $\hat{\bm{z}}$ represents the direction of the $d$-vector and $k_{\rm F}$ is the radius of the Fermi sphere. The function $\eta(\hat{k}_{z})$ has to be even $\eta(\hat{k}_{z})=\eta(-\hat{k}_{z})$, to satisfy the requirement ${\bm{d}}(\hat{\bm{k}})=-{\bm{d}}(-\hat{\bm{k}})$. For the chiral $p$-wave and $f$-wave pairing state, $\eta(\hat{k}_{z})$ is given by,
These $E_{1u}$ chiral order parameters generates two Weyl nodes at the north and south poles on the Fermi sphere and realizes WSCs with $\nu=1$. As Fig. 1 shows, the $E_{1u}$ chiral $f$-wave pairing involves the two horizontal line-nodes at $\hat{k}_{z}=\pm 1/\sqrt{5}$.
These chiral order parameters are relevant to candidate materials of WSCs. The chiral $p$-wave pairing, ${\bm{d}}(\hat{\bm{k}})=\Delta(\hat{k}_{x}\pm i\hat{k}_{y})\hat{\bm{z}}$, is established in the ABM state of the superfluid 3He under an ambient pressure [30, 31, 32, 33, 34, 35]. Although we consider the $D_{6h}$ symmetry, the calculated result with the chiral $p$-wave pairing in Eq. (2) is applicable to 3He because we do not consider the crystalline structure except for the form of the order parameter. The $E_{1u}$ chiral $f$-wave pairing is a candidate for the order parameter of UPt3 [46, 44].
Irreducible representation Basis function Chirality Parity
@@ II Model
Table 1: The irreducible representations, the basis functions, the chirality, the parity, and the existence of the extrinsic ATHE due to the short-range impurity potential for the chiral superconducting order for the point group $D_{6h}$.
@@ III Intrinsic anomalous thermal Hall effect
In chiral superconductors, there are intrinsic and extrinsic mechanisms for the ATHE. In this section, we present the expression of the intrinsic contribution to the THC and show its low temperature behavior is completely determined by the distribution of Weyl nodes in the momentum space.
For the model presented in Sec. II, two Weyl nodes are on the $k_{z}$ axis and these are separated by $\delta k_{\rm W}=2k_{\rm F}$ in the momentum space. When the WSCs are regarded as a family of two-dimensional superconductors labeled by $k_{z}$, their topological nature is revealed. Each subsystem labeled by $k_{z}$ is equivalent to a two-dimensional chiral superconductor and characterized by the Chern number defined as,
@@
where $E_{n}(\bm{k})$ is the quasiparticle energy in band $n$, and
@@
is the Berry curvature. In the case of WSCs described by Eq. (1) with the positive effective mass of the normal state, the momentum-dependent Chern number becomes, [71, 72]
@@
where $\Theta(x)$ is the Heaviside step function. The factor 2 in Eq. (5) arises from the spin degrees of freedom.
@@
Equation (7) clarifies that, at low temperatures, the intrinsic ATHE is completely determined by the distribution of Weyl nodes in the momentum space [16, 17, 18, 19]. Note that this low temperature formula is independent of $\eta(\hat{k}_{z})$ and thus the intrinsic ATHE is insensitive to additional line nodal excitations in WSCs at least to leading order in $T/T_{c}$.
To see that the intrinsic ATHE vanishes in the standard quasiclassical limit and does not appear in the Eilenberger framework [10], let us scale the thermal conductivity by $N(\epsilon_{\rm F})v_{\rm F}^{2}$. We find
@@
Equation (8) shows that the intrinsic contribution appears only at the first order in $1/(k_{\rm F}\xi_{0})$, whereas the the standard quasiclassical theory keeps only the terms of leading (zeroth) order in $1/(k_{\rm F}\xi_{0})$ [see Eq. (26)]. Hence, the intrinsic contribution to the ATHE drops out in the quasiclassical Eilenberger formalism  [77, 78], and we use the results above to compare with the extrinsic contribution to THC.
However, this hierarchy does not mean that the extrinsic contribution is always dominant. Recall that physically, the intrinsic ATHE is due to the gapless surface Majorana modes, which always give a $T$-linear contributions at low temperature, regardless of the existence of nodal excitations in bulk [6, 73, 79]. When the intrinsic contribution is derived from the bulk Hamiltonian using transport theory, the effect of the surface Majorana mode is incorporated into the THC via the correction to the Kubo formula due to the thermal magnetization current [74, 7, 75]. In contrast, whether the extrinsic ATHE exhibits $T$-linear behavior in the low temperature range, and, if it does, how large this contribution is, depends on the existence of nodal excitations and the location of impurity bands near the Fermi energy, see Ref. [22] and our analysis below. Therefore, whether intrinsic or extrinsic contributions dominate at low $T$ depends on the specifics of the material.

### v2

@@ Abstract
We investigate the anomalous thermal Hall effect (ATHE) in Weyl superconductors realized by the $E_{1u}$ ($p$-wave and $f$-wave) chiral superconducting order for the point group $D_{6h}$. Using the quasiclassical transport theory, we analyze the influence of the impurity scatterings and the line-nodal excitations on the ATHE. We compare the extrinsic (impurity-induced) ATHE with the intrinsic (topological) ATHE to identify the dominant contribution. Because the transverse response is sensitive to the slope in the density of states (DOS) at the Fermi energy, the extrinsic ATHE vanishes in both the Born (weak impurity potential) and unitarity (strong impurity potential) limits. The amplitude of the impurity contribution to the thermal Hall conductivity (THC) reaches maximum between these limits when the slope of the DOS becomes large due to impurity bands near the Fermi energy. In such situations, the extrinsic ATHE dominates the intrinsic ATHE even at low temperatures. The extrinsic ATHE is sensitive to line-nodal excitations, whereas the intrinsic ATHE is insensitive to bulk excitations. When line nodes involve the sign change of the order parameter, the impurity contribution to the THC is suppressed even though the phase space for low-energy excitation is large. In contrast, if line nodes are not accompanied by such sign changes, the extrinsic ATHE is significantly enhanced. Our results form a basis for the comprehensive analysis of anomalous thermal transport in Weyl superconductors.
@@ I Introduction
Weyl superconductors (WSCs) are time-reversal symmetry broken (TRSB) superconductors whose low-energy excitations behave as Weyl quasiparticles [1]. They are realized when the superconducting condensate consists of Cooper pairs with a fixed orbital angular momentum and are described by a complex order parameter, $\Delta(\bm{k})\propto(k_{x}\pm ik_{y})^{\nu}\;(\nu\in\mathbb{Z})$. On a three-dimensional Fermi surface centered on the $\Gamma$ point, the gap closing points at $k_{x}=k_{y}=0$ become Weyl points, which are sources and drains of Berry flux [2, 3, 4]. As a result, the near-nodal Bogoliubov quasiparticles behave as Weyl particles. It is common to both refer to such order parameters and label the corresponding ground states as chiral [5]. Generically, chiral superconductors with a three-dimensional Fermi surface are good candidates for the realization of WSCs.
@@
Among the phenomena that experimentally identify WSCs are the anomalous thermal Hall effect (ATHE), the transverse thermal current driven by a temperature gradient without an applied magnetic field, and the chiral anomaly-induced phenomena, such as the torsional chiral magnetic effect and the negative thermal magnetoresistivity by textures of the order parameters [6, 7, 8, 9, 10, 11, 12, 13]. In particular, the observation of the ATHE unambiguously identifies the chiral ground states.
@@
At the microscopic level, there are two sources for a finite thermal Hall signal in chiral superconductors: intrinsic and extrinsic. The intrinsic mechanisms arise from the effective Lorentz force generated by the Berry curvature and are characterized by the configuration of Weyl points in momentum space [14, 15, 16, 17, 18, 19]. The extrinsic mechanism relies on impurities and the transfer of the angular momentum between the condensate and the quasiparticles during scattering events [20, 21, 22, 23, 24], which contributes both to the energy-dependent skew scattering and the Andreev (inter-branch, electrons to holes and vice versa) scattering at impurity sites. The skew scattering directly couples to the temperature gradient for $|\nu|=1$ and results in the ATHE signal [20, 21, 22]. The Andreev mechanism only couples to impurities if the impurity potential is non-$s$-wave, with scattering matrix elements coupling different angular-momentum channels, and therefore appears for finite size impurities, dominating the response for $|\nu|\geq 2$ [23, 24, 25].
@@
In this paper, we address two aspects of the ATHE in chiral superconductors with point-like impurities. First, we elucidate its origin and demonstrate the relation between the extrinsic ATHE and the evolution of the density of states (DOS) in the sub-gap impurity band, which arises from the broadening of impurity resonant states [26, 27]. In particular, the particle-hole anisotropy, which is necessary for transverse transport, sensitively depends on the scattering phase-shift at individual impurities [28, 29]. Second, we focus on the magnitude of the thermal Hall conductivity (THC) in the situation where the winding number, $\nu$, differs from the total angular momentum of the Cooper pairs, $l$. Such a situation often occurs when the chiral superconducting order involves additional nodal excitations.
@@
Our results are relevant to several candidates of WSCs. Historically, the Anderson-Brinkmann-Morel (ABM) state in 3He was the first well-established Weyl superfluid with the chiral $p$-wave pairing [30, 1, 31, 32, 33, 34, 35]. In superconducting materials, the chiral ground states over three-dimensional Fermi surfaces were proposed in a number of materials, including URu2Si2, UPt3, U1-xThxBe13, and SrPtAs [36, 37, 38, 15, 39, 40, 41, 42, 43, 44, 45, 46, 47, 16, 48, 49, 50, 51, 52, 53, 54, 55]. The chiral $d$-wave pairing, $\Delta(\bm{k})\propto k_{z}(k_{x}\pm ik_{y})$, was proposed in URu2Si2 and explains the giant Nernst effect above the critical temperature due to preformed chiral Cooper pairs [39, 40]. Uranium compounds UPt3 and U1-xThxBe13 show spin-triplet superconductivity and multiple superconducting phases as a function of tuning parameters [56, 57, 58, 59, 60]. While the symmetry of the superconducting order parameter is still a subject of debate in these materials, TRSB was observed in the so-called B-phase, suggesting a chiral nature of the condensate [43, 61]. In general, ferromagnetic superconductors show a complex “nonunitary” order parameter and support Weyl nodes on three-dimensional Fermi surfaces [62, 63, 64, 65, 29]. Consequently, ferromagnetic UCoGe, URhGe, and UGe2 are also candidates of WSCs [66, 67, 68, 69].
@@
We employ the quasiclassical transport (Eilenberger) theory, which is a hierarchical expansion in $(k_{\rm F}\xi_{0})^{-1}\sim T_{\rm c}/T_{\rm F}\ll 1$, where $k_{\rm F}$ is the Fermi momentum, $\xi_{0}=v_{\rm F}/(2\pi T_{\rm c})$ is the superconducting coherence length, $v_{\rm F}$ is the Fermi velocity, and $T_{\rm c}$ and $T_{\rm F}$ are the superconducting transition and Fermi temperatures, respectively. Throughout this paper, we set $k_{\rm B}=\hbar=1$ for simplicity. In this theory, the extrinsic contribution is of leading order, while the intrinsic mechanism is smaller by a factor of $(k_{\rm F}\xi_{0})^{-1}$. Hence, understanding the extrinsic contribution is essential for correctly interpreting future experiments measuring the ATHE. While the intrinsic ATHE has been widely investigated in the context of topological material science [70, 71, 72, 73, 74, 7, 75, 16, 17, 18, 19], studies of the extrinsic ATHE have been very limited. Our work thus fills this gap for WSCs.
@@
The rest of this paper is organized as follows. We introduce the model of WSCs in Sec. II and present the expression for the intrinsic contribution to the THC in Sec. III. In Sec. IV, we review the quasiclassical transport theory, and in Sec. IV.3, we calculate the nonequilibrium quasiclassical Green function (QGF) under temperature gradients to obtain the thermal conductivity. The results of the low-temperature expansion analysis of the thermal conductivity are given in Sec. V. In Sec. VI, we associate the extrinsic ATHE with the formation of the impurity bands near the Fermi energy. In Sec. VII, we discuss how line-nodal excitations influence the ATHE, focusing on the opposing effects that nodal excitations have on longitudinal and anomalous Hall transport. Section VIII provides a summary and conclusion. Detailed derivations of several mathematically complex equations in the main text are given in the Appendices.
@@
Figure 1: Superconducting gap structures in the $E_{1u}$ chiral superconducting states considered in the main text [see Eqs. (1) and (2)]. The left and right panels illustrate the $E_{1u}$ chiral $p$-wave and $f$-wave states, respectively. The color of the superconducting gap represents the phase of the complex order parameter. The gray sphere indicates the underlying Fermi sphere.
Irreducible representation Basis function Winding number ($\nu$) Parity
@@
Table 1: The irreducible representations, the basis functions, the winding number ($\nu$), the parity, and the existence of the extrinsic ATHE due to the short-range impurity potential for the chiral superconducting order in the point group $D_{6h}$.
II Model
In this paper, we consider the chiral superconducting orders for the point group $D_{6h}$. Table. 1 summarizes the irreducible representations and their basis functions exhibiting the chiral superconducting order [76]. Among them, the skew scattering does not couple to the $E_{1g}$, $E_{2g}$ and $E_{2u}$ chiral superconducting order with the $\delta$-function type short-range impurity potential [23, 24]. As described in Sec. I, the $E_{2g}$ and $E_{2u}$ chiral superconducting orders have $|\nu|=2$, so that the skew scattering does not couple to these chiral orders. The $E_{1g}$ chiral superconducting order has $|\nu|=1$, but the line-node at the equator ($k_{z}=0$) leads to a precise cancellation of the skew scattering contribution to the ATHE [see Sec. V]. When impurities have finite size, $R\gtrsim 1/k_{\rm F}$, the Andreev scattering couples to these chiral superconducting orders and contributes to the ATHE. Hence, the extrinsic mechanism is negligible in the $E_{1g}$, $E_{2g}$, and $E_{2u}$ chiral states when the impurity potential is short-range.
The $E_{1u}$ chiral superconducting order includes both $p$-wave and $f$-wave pairings. As discussed below, these pairing states couple to the skew scattering, leading to the extrinsic ATHE. In this paper, we examine the chiral $p$-wave and $f$-wave pairings in the $E_{1u}$ state on a spherical Fermi surface to evaluate the effects of impurity bands and nodal excitations on the skew scattering. In these odd parity states, the $d$-vector is expressed by,
with the normalized crystal momentum $\hat{\bm{k}}={\bm{k}}/k_{\rm F}=(\hat{k}_{x},\hat{k}_{y},\hat{k}_{z})$. Here, $\hat{\bm{z}}$ represents the direction of the $d$-vector and $k_{\rm F}$ is the radius of the Fermi sphere. The function $\eta(\hat{k}_{z})$ has to be even $\eta(\hat{k}_{z})=\eta(-\hat{k}_{z})$, to satisfy the requirement ${\bm{d}}(\hat{\bm{k}})=-{\bm{d}}(-\hat{\bm{k}})$. For the $E_{1u}$ chiral $p$-wave and $f$-wave pairing states, $\eta(\hat{k}_{z})$ is given by,
These $E_{1u}$ chiral order parameters generate two Weyl points at the north and south poles on the Fermi sphere and realize WSCs with $\nu=1$. As Fig. 1 shows, the $E_{1u}$ chiral $f$-wave pairing involves, in addition, the two horizontal line nodes at $\hat{k}_{z}=\pm 1/\sqrt{5}$.
These chiral order parameters are relevant to candidate materials of WSCs. The chiral $p$-wave pairing, ${\bm{d}}(\hat{\bm{k}})=\Delta(\hat{k}_{x}\pm i\hat{k}_{y})\hat{\bm{z}}$, is established in the ABM state of the superfluid 3He under an ambient pressure [30, 31, 32, 33, 34, 35]. Although we consider the $D_{6h}$ point group, the calculated result with the chiral $p$-wave pairing is applicable to 3He because we do not consider the crystalline structure except for the form of the order parameter. The $E_{1u}$ chiral $f$-wave pairing has been discussed as a candidate for the order parameter of UPt3 [46, 44].
@@ III Intrinsic anomalous thermal Hall effect
In chiral superconductors, there are both intrinsic and extrinsic contributions to the ATHE. One of our goals is to compare the two contributions to identify the dominant contribution. In Sec. III, we present the expression for the intrinsic contribution to the THC and show that its low-temperature behavior is completely determined by the configuration of Weyl points in momentum space.
For the model presented in Sec. II, two Weyl points lie on the $k_{z}$ axis, separated by $\delta k_{\rm W}=2k_{\rm F}$ in momentum space. Let us regard the WSCs as a family of two-dimensional superconductors labeled by $k_{z}$ in order to understand their topological nature. Each subsystem labeled by $k_{z}$ is equivalent to a two-dimensional chiral superconductor and characterized by the $k_{z}$-resolved Chern number,
@@
where $E_{n}(\bm{k})$ is the quasiparticle energy and
@@
is the Berry curvature for the $n$-th band [2]. When the effective mass is positive in the normal state, the $k_{z}$-resolved Chern number in the WSCs described by Eq. (1) is given by,
@@
where $\Theta(x)$ is the Heaviside step function [71, 72]. The factor 2 in Eq. (5) arises from the spin degrees of freedom.
@@
Equation (7) clarifies that, at low temperatures, the intrinsic ATHE is completely determined by the configuration of Weyl points in momentum space [16, 17, 18, 19]. We note that this low-temperature formula does not depend on $\eta(\hat{k}_{z})$, indicating that the intrinsic ATHE is insensitive to additional line-nodal excitations in WSCs at least to leading order in $T/T_{c}$.
The intrinsic ATHE is the first order in $(k_{\rm F}\xi_{0})^{-1}$ and is not included in the quasiclassical transport (Eilenberger) theory [10]. To illustrate this, we scale the thermal conductivity by $N(\epsilon_{\rm F})v_{\rm F}^{2}$ and, for our choice of $\delta k_{\rm W}=2k_{\rm F}$, we find,
@@
Equation (8) shows that the intrinsic contribution appears only at the first order in $(k_{\rm F}\xi_{0})^{-1}$, whereas the the standard quasiclassical transport theory keeps only the terms of leading (zeroth) order in $(k_{\rm F}\xi_{0})^{-1}$ [see Eq. (28)]. Hence, the intrinsic contribution to the ATHE drops out in the quasiclassical Eilenberger formalism [77, 78].
However, this hierarchy does not mean that the extrinsic contribution is always dominant. Note that, physically, the gapless surface Majorana modes are responsible for the intrinsic ATHE and always bring about a $T$-linear contribution in the THC at low temperature, regardless of the existence of nodal excitations in bulk [6, 73, 79]. When the intrinsic contribution is derived from the bulk Hamiltonian using the Kubo theory, the effects of the surface Majorana mode are incorporated into the THC via a correction to the Kubo formula arising from the local equilibrium thermal magnetization current [74, 7, 75]. In contrast, whether the extrinsic ATHE exhibits $T$-linear behavior in the low-temperature range, and, if it does, how large this contribution is, depends on the existence of nodal excitations and the details of the impurity scattering [see Ref. [22] and our analysis below]. Therefore, whether intrinsic or extrinsic contributions dominate at the low temperature depends on the specifics of the material.